OpenStudy (bananas1234):

Use a trigonometric ratio to find the specified length to the nearest hundredth for each triangle. Show all work! X = 28.6 feet Find Y

OpenStudy (bananas1234):

 The triangle looks like this

OpenStudy (bananas1234):

@Directrix

OpenStudy (anonymous):

\[\frac{ y }{ x }=\cos 60,y=x \cos 60=28.6*\frac{ 1 }{ 2 }=?\]

OpenStudy (bananas1234):

it would be 14.3?

OpenStudy (anonymous):

14.30,though you are correct.

OpenStudy (bananas1234):

ok